Find a 2Γ—2 matrix A such that A^2 is a diagonal but not A


1
Expert's answer
2022-01-10T13:08:50-0500

Let

"A=\begin{bmatrix}
 a & b \\
 c & d
\end{bmatrix}"

Then


"A^2=\begin{bmatrix}
 a & b \\
 c & d
\end{bmatrix}\begin{bmatrix}
 a & b \\
 c & d
\end{bmatrix}=\begin{bmatrix}
 a^2+bc & ab+bd \\
 ac+dc & bc+d^2
\end{bmatrix}"

Given "A^2" is a diagonal. Then


"\begin{cases}
 ab+bd=0 \\
 ac+dc=0
\end{cases}=>\begin{cases}
 b(a+d)=0 \\
 c(a+d)=0
\end{cases}"

If "a+d\not=0," then "b=c=0," and matrix "A=\begin{bmatrix}
 a & 0 \\
 0 & d
\end{bmatrix}" is a diagonal.

We have contradiction.

Hence "a+d=0" and at least one of the numbers "c" and "b" is not equal to zero.

Therefore

"A=\begin{bmatrix}
 a & b \\
 c & -a
\end{bmatrix}," and at least one of the numbers "c" and "b" is not equal to zero.